# The Key Physics of Ice Premelting

### v1

@@ VII.1 Healing distance of a surface
Before embarking into the details, however, it is worth mentioning the concept of healing distance, that is important in the theory of wetting on patterned substrates.de Gennes, Brochard-Wyart, and Quéré (2004) The question here is, whether a thin film deposited on a patterned substrate will ’heal’ the holes left by the pattern and lead to a fully flat liquid-vapor surface, or, alternatively, whether it will adapt to the substrate and lead to a liquid-vapor surface that reveals the shape of the pattern. The decision depends on a balance between the free energy cost of increasing the surface area by following the pattern; and the free energy cost of healing the surface by changing the preferred equilibrium film thickness. The former cost is governed by the liquid-vapor surface tension, $\gamma_{{lv}}$, while the latter is controlled by the interface potential, $g(h_{e})$. The result emerging from the theory of wetting is that there is a relevant healing distance, or correlation length, $\xi=\sqrt{\gamma_{lv}/g^{\prime\prime}(h_{e})}$ (with $g^{\prime\prime}$ the second derivative of $g(h)$ with respect to $h$) which dictates the film’s decision. If the surface tension is very large, $\xi$ is large, and the film will heal all patterns of much smaller pitch to save the cost of creating extra surface area (as the films height remains correlated over length scales larger than the pattern relief); if, however, the interface potential is strong, $\xi$ is small, and the liquid film will follow any pattern of larger pitch, to save the free energy cost of perturbing the film thickness away from its equilibrium value (as the pattern relief is now of greater scale than the film height correlation length).
@@
Using our data for $g^{\prime\prime}$ reported in Fig.7 together with the surface tension of liquid water, we find that $\xi\approx 1$ nm for the basal face approaching $T_{t}$, which is somewhat larger (but in the order of magnitude) than the scale of a step, ca. one molecular diameter, and much smaller than the typical distance between steps observed in experiment. This suggests that, indeed, a premelting film could mimic faithfully the shape of the solid ice underneath, so that the apparent conflicting observations of ice premelting and surface steps can be reconciled.
@@ VII.2 Smooth and rough interfaces
Since both roughening and premelting are surface disorder phenomena, the question then is, whether the roughening transition must necessarily occur before the appearance of significant surface premelting or not. If this were the case, the observation of steps on the basal plane would imply there can be no premelting indeed, as suggested by Sazaki et al.Asakawa et al. (2015, 2016); Murata et al. (2016) But simple considerations suggest this need not be the case. Indeed, at the ice-water interface, the basal plane is thought to remain smooth up to the triple point. This expectation is supported by evidences of two dimensional nucleation of ice from its melt,Furukawa et al. (2021) the visual observation of elementary steps in vicinal planes,Murata et al. (2022) and the indication of layer-wise growth in molecular dynamics simulations.Nada and Furukawa (2005); Rozmanov and Kusalik (2012); Mitsui and Aoki (2023) Since the ice-water interface can be considered as a limiting case of a premelted surface with a film of infinite thickness, one then can expect intuitively that a buried basal surface under a thin premelting film should also be able to remain smooth up to the triple point.
@@
For a simple solid-liquid interface, the distinction between rough and smooth interfaces can be made quantitatively by the calculation of the surface height fluctuations. But for practical matters, it proofs convenient to better consider the problem by transforming the surface profile, $z({\bf x})$ into Fourier space, as $z({\bf q})$, with $\bf q$} a wave-vector along the plane of the interface. This is convenient, because it allows to obtain simple analytcal results for the Sine-Gordon Hamiltonian. Particularly, the mean squared surface fluctuations are given in Fourier space as:Chaikin and Lubensky (1995); Nelson, Piran, and Weinberg (2004)
@@
VII.3 Disgression: Can ice crystals be overheated?
@@ VIII The equilibrium structure of premelting films
Figure 12: Equilibrium structure of the ice basal face as described by the TIP4P/Ice model. (a) The probability density of film heights reveals two surface phase transitions as shown by the transformation from unimodal to bimodal and back to unimodal distributions as temperature increases. (b) snpashots of surface coverage of liquid-like molecules on top of the bulk ice. (c) Heat map of height distributions shows the emergence of thick correlated up and down domains at intermediate temperatures. (d) The spectrum of surface height fluctuations can be fitted to the an analytical model of Eq.(23) in order to estimate step free energies (c.f. Fig.13). Reproduced from Llombart et al.Llombart, Noya, and MacDowell (2020) Sci. Adv. 6 eaay9322 (2020). Distributed under a Creative Commons Attribution License $C$-BY.
@@
As an example, Fig.12-(a) shows the probability distribution of $\delta z_{{sl}}({\bf x})$ and $\delta z_{{lv}}({\bf x})$ as temperature increases from $\Delta T=T-T_{t}=-82$ K, to $\Delta T=-2$ K on the equilibrium basal surface of ice as described by the TIP4P/Ice model. At the lowest temperature, the distribution of surface heights is a narrow Gaussian, as expected for a sharp low temperature interface. In the language of the statistical thermodynamics of soft solid interfaces,Rommelse and den Nijs (1987); Prestipino, Santoro, and Tosatti (1995); Weichman and Prasad (1996); Woodraska and Jaszczak (1997); Jagla, Prestipino, and Tosatti (1999) this state is designed as an Ordered Flat phase (OF). As temperature increases, steps and terraces start to proliferate on the ice surface, as revealed by the multi-modal character of the probability distribution. A surface structure with proliferation of terraces but limited surface height fluctuations is known as a Disordered Flat Phase, as first identified by Rommelse and den Nijs in simple column models of the solid interface.Rommelse and den Nijs (1987) Finally, at yet higher temperature, the multi-modal probability distribution vanishes, and now becomes a broad Gaussian distribution. This transformation occurs concurrently with a significant increase of the premelting thickness, so it appears that the interface trades the entropy of step proliferation in the DOF phase, for the translational entropy of the liquid like water molecules. This last surface phase was named as a High-Temperature Reconstructed phase (HT-RF) in Ref.Llombart, Noya, and MacDowell, 2020, although the name of Flat Premelted phase (FP) would have been more appropriate.
@@
The study of the prism facet reveals also the existence of three different surface phases, albeit with the phase transitions occurring at significantly different temperatures from those observed for the basal face.Llombart, Noya, and MacDowell (2020) A related study for the TIP4P/2005 water model studied the height-height surface correlation in real space and also found evidence of the existence of three different phases on basal and prism faces of ice.Berrens, Bononi, and Donadio (2022)
@@ IX Premelting films away from the sublimation line
In practice, the ice-vapor interface can only be realized in equilibrium conditions for state points along the sublimation line, a situation which is actually more the exception than the rule. Actually, ice samples in the atmosphere will be most often found in conditions away from equilibrium.Libbrecht (2022) In this situation, the interface does no longer remain uniform, but will display pattern formation, inhomogeneities and instabilities.Misbah, Pierre-Louis, and Saito (2010) However, contrary to some simple situations, the departure from equilibrium cannot be fully interpreted just as a measure of the distance away from the sublimation line. The vicinity of the (metastable) liquid-vapor coexistence line means that, together with the sublimation/desublimation processes, it is necessary to account for the competing evaporation/condensation process of the marginally metastable liquid phase (c.f. Figure 4).
@@
However, the experiments performed by the Hokaido team show that water droplets on the ice surface form only in conditions of significant supersaturation over water vapor, when the pressure is increased beyond a temperature dependent threshold that lies well above the condensation line (c.f. Fig.).Asakawa et al. (2016); Murata et al. (2016)
@@
Particularly, it is known that the condition for mechanical equilibrium of a liquid droplet on an inert substrate is a positive Laplace pressure, $\Delta p>0$. So, likewise, in the quasi-stationary regime, with a living substrate feeding from the premelting film, the condition of mechanical equilibrium of the droplet now becomes $\Delta p_{k}>0$.Sibley et al. (2021) Solving for the marginal condition $\Delta p_{k}(p_{v},T)=0$ provides a kinetic condensation line, which dictates the pressure at which a droplet can first start growing atop the liquid/vapor surface. This can be worked out easily from the above equation and leads to a kinetic phase boundary line of Clausius-Clapeyron type:
@@
where $p_{kc}$ stand for the kinetic condensation pressure, and $\Delta H_{kc}>\Delta H_{{lv}}$ is an effective enthalpy of vaporization. This is always larger than the equilibrium vaporization enthalpy, $\Delta H_{{lv}}$, so that the condition for a droplet condensing on top of the growing premelted ice surface lies always above the equilibrium condensation line, in full agreement with the observations of the Hokaido team.Asakawa et al. (2016); Murata et al. (2016)
Another relevant phase line corresponds to the solution of Eq.(25) in the special case that the steady state thickness $h_{\rm qs}$ attains the value corresponding to the wetting spinodal, i.e., where the second derivative of the interface potential $g(h)$ first vanishes. The kinetic spinodal line lies above the kinetic coexistence line, and describes the phase points where a quasi-stationary regime can no longer hold. In this situation, the freezing rate at the ice/water surface can no longer match the condensation rate at the water/vapor surface, so that the premelting film thickness starts growing without bounds. This corresponds to a situation of kinetic complete surface melting, where the ice sample is freezing underneath the liquid film, but the liquid film thickness is diverging due to vapor condensation at a faster pace.Sibley et al. (2021)
@@
The different growth regimes discussed in this section are sketched in qualitative manner in Figure 11(d). However, an important thing to notice is that the quasi-stationary growth regime described here, and the corresponding kinetic boundary lines, are not a universal feature of the substance. As recognized by Kuroda and Gonda long time ago,Kuroda and Gonda (1990) they crucially depend on the growth mechanism followed by the crystal. Depending on the crystal plane, initial conditions and sample history, this can be nucleated, spiral or linear growth. The linear growth discussed above, however, is the one leading to the simplest algebraic results and is used here for the sake of illustration. In practice, in out of equilibrium conditions there occur film instabilities, and the mean field free energy model of Eq.(24), corresponding to films of uniform thickness is not sufficient to describe the complex out of equilibrium dynamic that actually takes place. However, accounting for inhomogeneities using the SG-CW free energy model of Section VII.B allows to describe complex non-equilibrium phenomena in remarkable agreement with experimental findings (c.f.: Figure 10).Sibley et al. (2021)
@@ X Step free energies and the Nakaya diagram
Using results from the sine-Gordon model of crystalline interfaces,Bennett et al. (1981); Chaikin and Lubensky (1995) the free energy cost to create a step on the ice/water surface amounts to $\beta_{{sl}}=\frac{2b^{2}}{\pi^{2}}\sqrt{\gamma_{{sl}}w}$.Noziéres and Gallet (1987) Now, for an interface exhibiting premelting, it is also necessary to account for the cost or rising the water/vapor surface due to the elevation of the step in the crystal surface below. By analogy with the previous result, it is expected that this free energy cost will be given by $\beta_{{lv}}=\frac{2b^{2}}{\pi^{2}}\sqrt{\gamma_{{lv}}g^{\prime\prime}}$.Llombart, Noya, and MacDowell (2020) The total free energy cost of creating a step on the premelted surface must then be given as $\beta=\beta_{{sl}}+\beta_{{lv}}$.Llombart, Noya, and MacDowell (2020)
The rationale of these two equations may be understood as follows: the deviation of a surface from planarity bears a free energy cost related to the increase of surface area, which can be measured as $\propto\gamma(dz/dx)^{2}dx$, with the square derivative of the surface profile $z(x)$ a measure of the surface area increment.Noziéres and Gallet (1987); Chaikin and Lubensky (1995) On average, the surface height increases by an amount equal to the full step height $dz=b$ over a horizontal distance on the scale of the surface correlation length, $dx=\xi$. This is $\xi_{{sl}}=\sqrt{\gamma_{{sl}}/w}$ or $\xi_{{lv}}=\sqrt{\gamma/g^{\prime\prime}}$ for either the ice/water and water/vapor surfaces, respectively. Then, the calculation of $\gamma(b/\xi)^{2}\xi$, leads to the previous results up to numerical factors.
@@
The alternating behavior of crystal growth rates of ice from the vapor has been reported in a number of experiments over the years,Kobayashi (1967); Lamb and D.Scott (1972); Sei and Gonda (1989); Libbrecht and Rickerby (2013) but a molecular explanation of this unusual behavior has remained elusive. However, thanks to the atomic scale resolution of computer simulations, the development of reliable force fields for water,Abascal et al. (2005); Abascal and Vega (2005) and the formulation of a suitable theoretical framework for the analysis of computer simulations,Benet et al. (2016, 2019); Llombart, Noya, and MacDowell (2020) a consistent explanation appears to emerge. The ice surface structure is extremely complex, and cannot be explained in terms of the traditional crystal surface models based on the surface diffusion of isolated adatoms that is adequate for the description of high energy solids very far away from their triple point. Contrary to this venerable picture, the ice surface exhibits all the way from 200 K to its melting point a significant amount of self-adsorption which increases with temperature and reaches the nanometer thickness in the neighborhood of the triple point. In this process, the basal and prism surfaces undergo a number of structural rearrangements,Llombart et al. (2020); Berrens, Bononi, and Donadio (2022) with profound impact on the crystal growth dynamics. The result of these transformations is that the growth rates become higher or lower for either the basal or prism facets in alternating order. Whereas the detailed structure of the surface phases is different, the underlying reason for the alternating growth rates is very much as anticipated by Kuroda and Lacmann many years ago.Kuroda and Lacmann (1982)
Refer to caption Refer to caption Refer to caption
@@
Figure 14: Interfacial structure and surface fluctuations of ice close to the triple point. (a) and (b) show the spectrum of surface fluctuations for the basal and prism faces in the near vicinity of the triple point as described by the TIP4P/2005 model of water. Results are displayed as wave-vector dependent surface tensions of the ice/water (blue) and water/vapor (red) surfaces, together with the coupled correlations (green). Symbols are results from computer simulations, while lines are results from the model of Eq.23. The corresponding surface tensions for interfaces of bulk water with either bulk ice or bulk vapor are shown as empty symbols and look equal to those of the premelting film up to $q=2$ nm-1, as indicated by the black dashed line. (c) Figure compares denstiy profiles of the premelting film (cotinuous lines) with those of the ice/water and water/vapor interfaces (dashed lines). The red and blue full lines are density profiles of solid-like and liquid-like molecules of the premelting film, respectively. The violet and green dashed lines are densities of solid-like and liquid-like molecules from the ice/water and water/vapor interfaces, respectively. Reproduced with permission from from Benet et al.Benet et al. (2016), Phys. Rev. Let. 117 096101 (2016). Copyright ©2016 American Physical Society.

### v2

@@ VII.1 Healing distance of a surface
Before embarking into the details, however, it is worth digressing on the healing distance, a concept that is important in the theory of wetting on patterned substrates.de Gennes, Brochard-Wyart, and Quéré (2004) The question here is, whether a thin film deposited on a patterned substrate will ’heal’ the holes left by the pattern and lead to a fully flat liquid-vapor surface, or, alternatively, whether it will adapt to the substrate and lead to a liquid-vapor surface that reveals the shape of the pattern. The decision depends on a balance between the free energy cost of increasing the surface area by following the pattern; and the free energy cost of healing the surface by changing the preferred equilibrium film thickness. The former cost is governed by the liquid-vapor surface tension, $\gamma_{{lv}}$, while the latter is controlled by the interface potential, $g(h_{e})$. The result emerging from the theory of wetting is that there is a relevant healing distance, or correlation length, $\xi=\sqrt{\gamma_{lv}/g^{\prime\prime}(h_{e})}$ (with $g^{\prime\prime}$ the second derivative of $g(h)$ with respect to $h$) which dictates the film’s decision. If the surface tension is very large, $\xi$ is large, and the film fills all holes with lateral dimension smaller than the healing distance to save the cost of creating extra surface area (as the film’s height remains correlated over length scales larger than the pattern relief); if, however, the interface potential is strong, $\xi$ is small, and the liquid film will follow the shape of the pattern, to save the free energy cost of perturbing the film thickness away from its equilibrium value (as the pattern relief is now of greater scale than the film height correlation length).
@@
Using our data for $g^{\prime\prime}$ reported in Fig.7 together with the surface tension of liquid water, we find that $\xi\approx 1$ nm for the basal face approaching $T_{t}$. This is somewhat larger (but in the order of magnitude) than the scale of a step, ca. one molecular diameter, but much smaller than the typical distance between steps observed in experiment. This suggests that, indeed, a premelting film could mimic faithfully the shape of the solid ice underneath, so that the apparent conflicting observations of ice premelting and surface steps can be reconciled.
@@ VII.2 Smooth and rough interfaces
Since both roughening and premelting are surface disorder phenomena, the question then is, whether the roughening transition must necessarily occur before the appearance of significant surface premelting or not. If this were the case, the observation of steps on the basal plane would imply there can be no premelting indeed, as suggested by Sazaki et al.Asakawa et al. (2015, 2016); Murata et al. (2016) But simple considerations suggest this need not be the case. Indeed, at the ice-water interface, the basal plane is thought to remain smooth up to the triple point. This expectation is supported by evidences of two dimensional nucleation of ice from its melt,Furukawa et al. (2021) the visual observation of elementary steps in vicinal planes,Murata et al. (2022) and the indication of layer-wise growth in molecular dynamics simulations.Nada and Furukawa (2005); Rozmanov and Kusalik (2012); Mochizuki, Murata, and Zhang (2023) Since the ice-water interface can be considered as a limiting case of a premelted surface with a film of infinite thickness, one then can expect intuitively that a buried basal surface under a thin premelting film should also be able to remain smooth up to the triple point.
@@
For a simple solid-liquid interface, the distinction between rough and smooth interfaces can be made quantitatively by the calculation of the surface height fluctuations. But for practical matters, it proofs convenient to better consider the problem by transforming the surface profile, $z({\bf x})$ into Fourier space, as $z({\bf q})$, with $\bf q$} a wave-vector along the plane of the interface. This is convenient, because it allows to obtain simple analytical results for the Sine-Gordon Hamiltonian. Particularly, the mean squared surface fluctuations are given in Fourier space as:Chaikin and Lubensky (1995); Nelson, Piran, and Weinberg (2004)
@@
VII.3 Digression: Can ice crystals be overheated?
@@ VIII The equilibrium structure of premelting films
Figure 12: Equilibrium structure of the ice basal face as described by the TIP4P/Ice model. (a) The probability density of film heights reveals two surface phase transitions as shown by the transformation from unimodal to bimodal and back to unimodal distributions as temperature increases. (b) snapshots of surface coverage of liquid-like molecules on top of the bulk ice. (c) Heat map of height distributions shows the emergence of thick correlated up and down domains at intermediate temperatures. (d) The spectrum of surface height fluctuations can be fitted to the an analytical model of Eq.(23) in order to estimate step free energies (c.f. Fig.13). Reproduced from Llombart et al.Llombart, Noya, and MacDowell (2020) Sci. Adv. 6 eaay9322 (2020). Distributed under a Creative Commons Attribution License $C$-BY.
@@
As an example, Fig.12-(a) shows the probability distribution of $\delta z_{{sl}}({\bf x})$ and $\delta z_{{lv}}({\bf x})$ as temperature increases from $\Delta T=T-T_{t}=-82$ K, to $\Delta T=-2$ K on the equilibrium basal surface of ice as described by the TIP4P/Ice model. At the lowest temperature, the distribution of surface heights is a narrow Gaussian, as expected for a sharp low temperature interface. In the language of the statistical thermodynamics of soft solid interfaces,Rommelse and den Nijs (1987); Prestipino, Santoro, and Tosatti (1995); Weichman and Prasad (1996); Woodraska and Jaszczak (1997); Jagla, Prestipino, and Tosatti (1999) this state is designed as an Ordered Flat phase (OF). As temperature increases, steps and terraces start to proliferate on the ice surface, as revealed by the multi-modal character of the probability distribution. A surface structure with proliferation of terraces but limited surface height fluctuations is known as a Disordered Flat Phase (DOF), as first identified by Rommelse and den Nijs in simple column models of the solid interface.Rommelse and den Nijs (1987) Finally, at yet higher temperature, the multi-modal probability distribution vanishes, and now becomes a broad Gaussian distribution. This transformation occurs concurrently with a significant increase of the premelting thickness, so it appears that the interface trades the entropy of step proliferation in the DOF phase, for the translational entropy of the liquid like water molecules. This last surface phase was named as a High-Temperature Reconstructed phase (HT-RF) in Ref.Llombart, Noya, and MacDowell, 2020, although the name of Flat Premelted phase (FP) would have been more appropriate.
@@
The study of the prism facet reveals also the existence of three different surface phases, albeit with the phase transitions occurring at significantly different temperatures from those observed for the basal face.Llombart, Noya, and MacDowell (2020) A related study for the TIP4P/2005 water model studied the height-height surface correlations in real space and also found evidence of the existence of three different phases on basal and prism faces of ice.Berrens, Bononi, and Donadio (2022)
@@ IX Premelting films away from the sublimation line
In practice, the ice-vapor interface can only be realized in equilibrium conditions for state points along the sublimation line, a situation which is actually more the exception than the rule. Indeed , ice samples in the atmosphere will be most often found in conditions away from equilibrium.Libbrecht (2022) In this situation, the interface does no longer remain uniform, but will display pattern formation, inhomogeneities and instabilities.Misbah, Pierre-Louis, and Saito (2010) However, contrary to some simple situations, the departure from equilibrium cannot be fully interpreted just as a measure of the distance of the vapor away from the sublimation line. The vicinity of the (metastable) liquid-vapor coexistence line means that, together with the sublimation/desublimation processes, it is necessary to account for the competing evaporation/condensation process of the marginally metastable liquid phase (c.f. Figure 4).
@@
However, the experiments performed by the Hokaido team show that water droplets on the ice surface form only in conditions of significant supersaturation over water vapor, when the pressure is increased beyond a temperature dependent threshold that lies well above the condensation line (c.f. Fig.11). Asakawa et al. (2016); Murata et al. (2016)
@@
Particularly, it is known that the condition for mechanical equilibrium of a liquid droplet on an inert substrate is a positive Laplace pressure, $\Delta p>0$. So, likewise, in the quasi-stationary regime, with a living substrate feeding from the premelting film, the condition of mechanical equilibrium of the droplet now becomes $\Delta p_{k}>0$.Sibley et al. (2021) Solving for the marginal condition $\Delta p_{k}(p_{v},T)=0$ provides a kinetic condensation line, which dictates the pressure at which a droplet can first start growing atop the liquid/vapor surface (c.f. dashed blue line of Fig.4) . This can be worked out easily from the above equation and leads to a kinetic phase boundary line of Clausius-Clapeyron type:
@@
where $p_{kc}$ stand for the kinetic condensation pressure, and $\Delta H_{kc}>\Delta H_{{lv}}$ is an effective enthalpy of vaporization. This is always larger than the equilibrium vaporization enthalpy, $\Delta H_{{lv}}$, so that the condition for a droplet condensing on top of the growing premelted ice surface lies always above the equilibrium condensation line, in full agreement with the observations of the Hokaido team (c.f. red line of Fig.11.E) .Asakawa et al. (2016); Murata et al. (2016)
Another relevant phase line corresponds to the solution of Eq.(25) in the special case that the steady state thickness $h_{\rm qs}$ attains the value corresponding to the wetting spinodal, i.e., where the second derivative of the interface potential $g(h)$ first vanishes. The kinetic spinodal line lies above the kinetic coexistence line, and describes the phase points where a quasi-stationary regime can no longer hold (c.f. dashed red line of Fig.4) . In this situation, the freezing rate at the ice/water surface can no longer match the condensation rate at the water/vapor surface, so that the premelting film thickness starts growing without bounds. This corresponds to a situation of kinetic complete surface melting, where the ice sample is freezing underneath the liquid film, but the liquid film thickness is diverging due to vapor condensation at a faster pace.Sibley et al. (2021)
@@
The different growth regimes discussed in this section are sketched in qualitative manner in Figure  4, to be compared with the experimental results of Fig. 11.E . However, an important thing to notice is that the quasi-stationary growth regime described here, and the corresponding kinetic boundary lines, are not a universal feature of the substance. As recognized by Kuroda and Gonda long time ago,Kuroda and Gonda (1990) they crucially depend on the growth mechanism followed by the crystal. Depending on the crystal plane, initial conditions and sample history, this can be nucleated, spiral or linear growth. The linear growth discussed above, however, is the one leading to the simplest algebraic results and is used here for the sake of illustration.
In practice, in out of equilibrium conditions there occur film instabilities, and the mean field free energy model of Eq.(24), corresponding to films of uniform thickness is not sufficient to describe the complex out of equilibrium dynamic that actually takes place. However, accounting for inhomogeneities using the SG-CW free energy model of Section VII.B allows to describe complex non-equilibrium phenomena in remarkable agreement with experimental findings (c.f.: Figure 10).Sibley et al. (2021)
@@ X Step free energies and the Nakaya diagram
Using results from the sine-Gordon model of crystalline interfaces,Bennett et al. (1981); Chaikin and Lubensky (1995) the free energy cost to create a step on the ice/water surface amounts to $\beta_{{sl}}=\frac{2b^{2}}{\pi^{2}}\sqrt{\gamma_{{sl}}w}$.Noziéres and Gallet (1987) In the presence of premelting it is also required to account for the cost of rising the water/vapor surface above the step so formed. By analogy with the previous result, it is expected that this free energy cost will be given by $\beta_{{lv}}=\frac{2b^{2}}{\pi^{2}}\sqrt{\gamma_{{lv}}g^{\prime\prime}}$.Llombart, Noya, and MacDowell (2020) The total free energy cost of creating a step on the premelted surface must then be given as $\beta=\beta_{{sl}}+\beta_{{lv}}$.Llombart, Noya, and MacDowell (2020)
The rationale of these two equations may be understood as follows: the deviation of a surface from planarity bears a free energy cost related to the increase of surface area, which can be measured as $\gamma(dz/dx)^{2}dx$, with the squared derivative of $z(x)$, used as a measure of the surface area increment.Noziéres and Gallet (1987); Chaikin and Lubensky (1995) On average, the surface height increases by an amount equal to the full step height $dz=b$ over a horizontal distance on the scale of the surface correlation length, $dx=\xi$. This is $\xi_{{sl}}=\sqrt{\gamma_{{sl}}/w}$ or $\xi_{{lv}}=\sqrt{\gamma/g^{\prime\prime}}$ for either the ice/water and water/vapor surfaces, respectively. Then, the calculation of $\gamma(b/\xi)^{2}\xi$, leads to the previous results up to numerical factors.
@@
The alternating behavior of crystal growth rates of ice from the vapor has been reported in a number of experiments over the years,Kobayashi (1967); Lamb and D.Scott (1972); Sei and Gonda (1989); Libbrecht and Rickerby (2013) but a molecular explanation of this unusual behavior has remained elusive. However, thanks to the atomic scale resolution of computer simulations, the development of reliable force fields for water,Abascal et al. (2005); Abascal and Vega (2005) and the formulation of a suitable theoretical framework for the analysis of computer simulations,Benet et al. (2016, 2019); Llombart, Noya, and MacDowell (2020) a consistent explanation appears to emerge. The ice surface structure is extremely complex, and cannot be explained in terms of the traditional crystal surface models based on the surface diffusion of isolated ad-atoms that is adequate for the description of high energy solids very far away from their triple point. Contrary to this venerable picture, the ice surface exhibits all the way from 200 K to its melting point a significant amount of self-adsorption which increases with temperature and reaches the nanometer thickness in the neighborhood of the triple point. In this process, the basal and prism surfaces undergo a number of structural rearrangements,Llombart et al. (2020); Berrens, Bononi, and Donadio (2022) with profound impact on the crystal growth dynamics. The result of these transformations is that the growth rates become higher or lower for either the basal or prism facets in alternating order. Whereas the detailed structure of the surface phases is different, the underlying reason for the alternating growth rates is very much as anticipated by Kuroda and Lacmann many years ago.Kuroda and Lacmann (1982)
[…]
@@
Figure 14: Interfacial structure and surface fluctuations of ice close to the triple point. (a) and (b) show the spectrum of surface fluctuations for the basal and prism faces in the near vicinity of the triple point as described by the TIP4P/2005 model of water. Results are displayed as wave-vector dependent surface tensions of the ice/water (blue) and water/vapor (red) surfaces, together with the coupled correlations (green). Symbols are results from computer simulations, while lines represent fits to the model of Eq.23. The corresponding surface tensions for interfaces of bulk water with either bulk ice or bulk vapor are shown as empty symbols and look equal to those of the premelting film up to $q=2$ nm-1, as indicated by the black dashed line. (c) Figure compares density profiles of the premelting film (continuous lines) with those of the ice/water and water/vapor interfaces (dashed lines). The red and blue full lines are density profiles of solid-like and liquid-like molecules of the premelting film, respectively. The violet and green dashed lines are densities of solid-like and liquid-like molecules from the ice/water and water/vapor interfaces, respectively. Reproduced with permission from from Benet et al.Benet et al. (2016), Phys. Rev. Let. 117 096101 (2016). Copyright ©2016 American Physical Society.